# The NTU Partitioned Matching Game for International Kidney Exchange Programs

### v1

@@ 2.3 Existence of a strong core matching
Our aim in this subsection is to prove that it can be decided in polynomial time if the strong core is non-empty, and if so, a matching in the strong core can be found in polynomial time. These results require a careful analysis of the structure of $M_{0}$-alternating paths, so the proof will feature several structural lemmas. We start by introducing some notation.
Let $\mathcal{SC}$ denote the set of matchings in the strong core. Let $\mathcal{K}$ denote the set of players for which there is no $M_{0}$-alternating cycle through the corresponding player edge. We denote by $\mathcal{K}^{*}$ the set of players $A$ for which $\exists M\in\mathcal{SC}$: $|V(M)\cap A|\leq 1$. Of course, we cannot determine $\mathcal{K}^{*}$ without deciding whether the strong core is empty, so we will not use $\mathcal{K}^{*}$ in the algorithm, only in the proof of correctness. By Lemma 5 point a), we have $\mathcal{K}^{*}\subseteq\mathcal{K}$.
@@
We introduce some additional notation and definitions related to the structure of $M_{0}$-alternating paths, as listed below.
  • An $M_{0}$-alternating $\delta$-path (or $\delta$-path for short) is an odd cycle and an $M_{0}$-alternating path with a common vertex $v$, where the $M_{0}$-edge adjacent to $v$ is on the path, and the cycle is $M_{0}$-alternating apart from $v$.
@@
  • An $(A,B;C)$-$\delta$-path for $A,B,C\in\mathcal{K}$ is an $M_{0}$-alternating $\delta$-path where $A,B$ are in the cycle and $C$ is the end of the path. The vertex of $C$ at the end of the path is called the tip. Note that the roles of $A$ and $B$ are exchangeable in the definition. See Figure 1 for an illustration.
  • The two vertices of a player $A$ are usually denoted by $a$ and $a^{\prime}$. For $A\in\mathcal{K}$ and a given $M\in\mathcal{SC}$, we usually denote the path in $M\Delta M_{0}$ containing $A$ by $P_{A}$, and the two endpoints of the path by $x_{A}$ and $y_{A}$. In some contexts, $P_{A}$ will be an arbitrary $M_{0}$-alternating path that contains $A$.
@@
A general proof technique is to consider a graph $G^{\prime}$ obtained from $E\cup M_{0}$ by removing a single player edge $A=aa^{\prime}\in M_{0}$, and adding some additional edges, so that $M_{0}\setminus\{A\}$ is a maximum matching in $G^{\prime}$. In such a case, $G^{\prime}$ has a Gallai-Edmonds decomposition with hypomatchable odd components $Q_{1},\dots,Q_{k+2}$, where $a\in Q_{k+1}$ and $a^{\prime}\in Q_{k+2}$. We denote the player edge entering $Q_{j}$ by $S_{j}=s_{j}s^{\prime}_{j}$ ($j\in[k]$), where $s^{\prime}_{j}\in Q_{j}$. We will frequently use the following property of alternating paths: if an $M_{0}$-alternating path starts in some $Q_{i}$ ($i\in[k+2])$, does not use $S_{i}$, and ends with an edge of $M_{0}$, then the endpoint is in some $Q_{j}$ ($j\in[k+2]$). Furthermore, if $j\neq i$, then $j\in[k]$, and the path enters $Q_{j}$ at the edge $S_{j}$. This also means that such a path cannot enter an odd component $Q_{j}$ twice. See Figure 2 for an illustration. We will also use the corollary that if $b\in Q_{j}$ for some $j\in[k+2]$ and there are $M_{0}$-alternating paths to $b$ from both $a$ and $a^{\prime}$, then $j\in[k]$.
Figure 2: An $M_{0}$-alternating path in $G^{\prime}$ from $Q_{i}$ to $Q_{j}$, indicated by thick edges. Edges in $M_{0}$ are solid, while edges in $E$ are dashed.
The following lemma enables us to obtain an alternating path from the union of two paths.
Lemma 7.
Let $P$ be an $M_{0}$-alternating path from player-edge $A\in\mathcal{K}$ to player-edge $B$, with endpoints $a$ and $b$, and let $P_{A}$ be an $M_{0}$-alternating path through $A$, with endpoints $x_{A}$ and $y_{A}$. Then there exists an $M_{0}$ alternating path $P^{\prime}\subseteq P\cup P_{A}$ from $b$ to either $x_{A}$ or to $y_{A}$ (with $M_{0}$-edges at both ends) that passes through $A$.
Proof.
We assume for contradiction that there is no such path. Define $G^{\prime}=(P\cup P_{A})\cup\{x_{A}b,y_{A}b\}\setminus\{A\}$. By this assumption, there is no $M_{0}$-alternating path between $a$ and $a^{\prime}$ in $G^{\prime}$, so we can consider the GE decomposition for $G^{\prime}$ (otherwise, as $A\in\mathcal{K}$, it contains $x_{A}b$ or $y_{A}b$ and gives such an $M_{0}$-alternating path in $G$). We may assume that $x_{A}$ is closer to $a$ than to $a^{\prime}$ on $P_{A}$. We show that the vertices $b,x_{A},y_{A}$ are in the same $Q_{i}$ for some $i\in[k]$. Indeed, there are $M_{0}$-alternating paths to all of them from $a$ or $a^{\prime}$ ending in $M_{0}$-edges, so they are in odd components of the GE decomposition, and they must be in the same $Q_{i}$ because of the edges $x_{A}b$ and $y_{A}b$. The $a-x_{A}$ path implies that $i\neq k+2$, and the $a^{\prime}-y_{A}$ path implies that $i\neq k+1$. Thus, the edge $S_{i}$ is contained in both the $a-x_{A}$ subpath and the $a^{\prime}-y_{A}$ subpath of $P_{A}$; but these two subpaths are disjoint, a contradiction. ∎
Note that the above proof works even if $y_{a}=b$; in this case, the path obtained is from $b$ to $x_{A}$. The following is an immediate corollary of Lemma 7.
Proposition 8.
Let $A,B\in\mathcal{K}$, and let $P$ be an $M_{0}$-alternating path between $A$ and $B$. Then there is no $M\in\mathcal{SC}$ such that $|V(M)\cap A|=2$ and $|V(M)\cap B|=0$.
Proof.
Suppose that $M$ is such a matching, and let $x_{A}$ and $y_{A}$ denote the two endpoints of the path $P_{A}$ containing $A$ in $M\Delta M_{0}$. Let $a$ and $b$ be the two endpoints of $P$. By Lemma 7, there exists an $M_{0}$-alternating path $P^{\prime}\subseteq P\cup P_{A}$ from $b$ to either $x_{A}$ or to $y_{A}$ (and to one that is different from $b$), that passes through $A$. But then $P^{\prime}$ is a blocking path for $M$, a contradiction. ∎
Lemma 9.
Consider an $(A,B;C)$-$\delta$-path $H$, and an $M\in\mathcal{SC}$ such that $|V(M)\cap A|=1$ and $|V(M)\cap C|=2$. Then there is an $(A,B;C^{\prime})$-$\delta$-path $H^{\prime}$ with the same cycle as $H$, such that the path of $H^{\prime}$ contains $C$, and $C^{\prime}$ is one end of the path in $M\Delta M_{0}$ containing $C$ (hence $C^{\prime}\in\mathcal{K^{*}})$.
Proof.
Let $P_{C}$ be the path in $M\Delta M_{0}$ that contains $C$, and let $x_{C}$ and $y_{C}$ be its two endpoints. Define $G^{\prime}$ by adding the edge $x_{C}y_{C}$ and removing $A$. Let $C=cc^{\prime}$, where $c$ is the tip of $H$. The $(A,B;C)$-$\delta$-path $H$ contains an $a-c$ path and an $a^{\prime}-c$ path, so $c\in Q_{i}$ for some $i\in[k]$. Let $P_{1}$ (respectively $P_{2}$) denote the $a-s_{i}$ part of the $a-c$ path ($a^{\prime}-s_{i}$ part of the $a^{\prime}-c$ path, respectively). The part from $s_{i}$ to $c$ is common; let this be denoted by $P$. We can use Lemma 7 for $P$ and $P_{C}$, to obtain a path $P^{\prime}\subseteq P\cup P_{C}$ from $s_{i}$ to $x_{C}$ or $y_{C}$ (we may assume $x_{C}$), that starts with $S_{i}$ and contains $C$. $P^{\prime}$ enters $Q_{i}$ only once (at its start), so it cannot intersect $P_{1}$ or $P_{2}$ before reaching $C$. It also cannot intersect them after reaching $C$ either, because that would create an $M_{0}$-alternating cycle through $C$. Let $C^{\prime}$ be the player that contains $x_{C}$. The union of $P_{1},P_{2}$, and $P^{\prime}$ gives the desired structure. ∎
It follows from the lemma that if there is an $(A,B;C)$-$\delta$-path and an $M\in\mathcal{SC}$ such that $|V(M)\cap A|=1$, then $|V(M)\cap B|=2$. Indeed, otherwise the $(A,B;C^{\prime})$-$\delta$-path given by the lemma would contain a blocking path for $M$ by Lemma 5.
Lemma 10.
Let $P$ be an $M_{0}$-alternating $A-B-C$ path for some $A\in\mathcal{K}$, $B\in\mathcal{K}^{*}$, $C\in\mathcal{K}$. Then there exists an $(A,B;C)$-$\delta$-path or a $(C,B;A)$-$\delta$-path.
Proof.
Let $a,a^{\prime},b,b^{\prime},c,c^{\prime}$ be the vertices of $A,B,C$ in the order in which they appear in $P$. Let $M\in\mathcal{SC}$ be a matching such that $|V(M)\cap B|=1$, and let $P_{A},P_{B},P_{C}$ be the paths in $M\Delta M_{0}$ that contain $A$, $B$ and $C$, respectively; note that these paths are not necessarily distinct. We assume that one endpoint of $P_{B}$ is $b$; let the other endpoint be denoted by $x_{B}$. Let $X\subseteq V$ denote the set of endpoints of $P_{A}$, $P_{B}$, and $P_{C}$, except for $b$.
We define $G^{\prime}$ by adding a clique on $X$ and removing $B$. If there is an $M_{0}$-alternating path in $G^{\prime}$ between $b$ and $b^{\prime}$, then it must contain a new edge as $B\in\mathcal{K}$ and as we have a clique on $X$, we can assume that it contains exactly one new edge. However, deleting this edge and adding back $B$ we get an $M_{0}$-alternating path in $G$ between two players with one vertex covered in $M$ that covers both vertices of $B$, contradicting $M\in\mathcal{SC}$ by Lemma 5. Since there is no alternating path between $b$ and $b^{\prime}$ in $G^{\prime}$, we can consider the GE-decomposition. The existence of $P_{B}$ implies that $x_{B}$ is in $Q_{i}$ for some $i\in[k+1]$, since $P_{B}$ cannot enter $Q_{k+2}$. As $P$ contains a $b-a$ subpath and a $b^{\prime}-c^{\prime}$ subpath that are disjoint, it is not possible for both $a$ and $c^{\prime}$ to be in $Q_{i}$. Indeed, if $i=k+1$, then $c^{\prime}\notin Q_{i}$, while if $i\in[k]$, then $S_{i}$ cannot be on both paths (but it is included in any $M_{0}$-alternating path from $\{b,b^{\prime}\}$ that enters $Q_{i}$).
Case 1: $a\in Q_{j}$ for some $j\neq i$, $c^{\prime}\in Q_{i}$.
In this case, $i\in[k]$ and $j\in[k+1]$. Because of the clique on $X$, one endpoint of $P_{A}$ must be in $Q_{i}$. If both endpoints are in $Q_{i}$, then there is a path to $Q_{i}$ that leaves $Q_{j}$ not on $S_{j}$, so it must enter $Q_{i}$ using $S_{i}$. If one endpoint is $b$, then the path from $b$ to $a$ must contain $S_{j}$ if it exists, so we get $S_{i}\in P_{A}$ again. As $S_{i}$ is on $P_{A}$ and is also on $P_{B}$, we have $P_{A}=P_{B}$.
Let $P_{1}$ be the part of $P$ from $b^{\prime}$ to $c^{\prime}$. Let $P_{2}$ be the part of $P$ from $b$ to $s_{j}$ if $j\in[k]$, and empty if $j=k+1$. Let $P_{3}$ be the part of $P_{B}$ from $s_{j}$ (or from $b$ if $j=k+1$) to the first intersection with $P_{1}$. Such an intersection exists because $S_{i}\in P_{1}\cap P_{B}$. Note also that $P_{3}$ traverses $A$ (as $P_{B}$ cannot enter $Q_{j}$ again and $A\in Q_{j}$) and it can intersect $P_{2}$ only in $s_{j}$, because another intersection would give an $M_{0}$-alternating cycle through $A$. Thus, $P_{1}\cup P_{2}\cup P_{3}\cup\{B\}$ is an $(A,B;C)$-$\delta$-path.
Case 2: $c^{\prime}\in Q_{j}$ for some $j\neq i$, $a\in Q_{i}$.
In this case, $i\in[k+1]$ and $j\in[k]\cup\{k+2\}$. Because of the clique on $X$, one endpoint of $P_{C}$ must be in $Q_{i}$. If both endpoints are in $Q_{i}$, then there is a path to $Q_{i}$ that leaves $Q_{j}$ not on $S_{j}$, so it must enter $Q_{i}$ using $S_{i}$, implying $P_{B}=P_{C}$. If one endpoint is $b$, then $P_{B}=P_{C}$ is obvious. We also get $j\neq k+2$.
We define $P_{1}$ as the part of $P$ from $b$ to $a$, $P_{2}$ to be the part of $P$ from $b^{\prime}$ to $s_{j}$, and $P_{3}$ to be the part of $P_{C}$ from $s_{j}$ to the first intersection with $P_{1}$. By a similar argument as in Case 1, $P_{1}\cup P_{2}\cup P_{3}\cup\{B\}$ is a $(C,B;A)$-$\delta$-path.
Case 3: $a\in Q_{j}$ for some $j\neq i$, $c^{\prime}\in Q_{\ell}$ for some $\ell\neq i$.
Because of the clique on $X$, $S_{i}\in P_{A}\cap P_{B}\cap P_{C}$, so $i\in[k]$ and $P_{A}=P_{B}=P_{C}$, which also implies $\ell\in[k]$. If the order of these player-edges on $P_{B}$ is $B,A,C$, then we define $P_{1},P_{2},P_{3}$ as in Case 1. The argument is almost the same as Case 1, except that $P_{3}$ intersects $P_{1}$ because $S_{\ell}\in P_{1}\cap P_{B}$. As in Case 1, $P_{1}\cup P_{2}\cup P_{3}\cup\{B\}$ is an $(A,B;C)$-$\delta$-path.
If the order of the player-edges in $P_{B}$ is $B,C,A$, then we define $P_{1}$ as the part of $P$ from $b$ to $a$, $P_{2}$ as the part of $P$ from $b^{\prime}$ to $s_{\ell}$, and $P_{3}$ as the part of $P_{B}$ from $s_{\ell}$ to the first intersection with $P_{1}$. $P_{3}$ intersects $P_{1}$ because $S_{j}\in P_{1}\cap P_{B}$. As in Case 2, $P_{1}\cup P_{2}\cup P_{3}\cup\{B\}$ is a $(C,B;A)$-$\delta$-path. ∎
Proposition 11.
Let $A,B,C\in\mathcal{K}$ and consider an $(A,B;C)$-$\delta$-path, and an $M\in\mathcal{SC}$ such that $|V(M)\cap A|=1$ and $|V(M)\cap B|=2$. Then $A$ and $B$ are on the same alternating path in $M\Delta M_{0}$.
Proof.
First, we consider the case $|V(M)\cap C|=1$. Suppose for contradiction that $A$ and $B$ are on different alternating paths in $M\Delta M_{0}$: $P_{A}$ with other endpoint $x_{A}$, and $P_{B}$ with endpoints $x_{B}$ and $y_{B}$. Let $c^{\prime}$ be the tip of the $\delta$-path. We define $G^{\prime}$ by adding a complete graph on $x_{A},x_{B},y_{B},c^{\prime}$ and removing $A$. Then there is no $M_{0}$-alternating path between $a$ and $a^{\prime}$, so we can consider the GE-decomposition (otherwise we get an alternating cycle through $A$ or a weakly blocking coalition).
The vertices $x_{A}$ and $c^{\prime}$ are in the same $Q_{i}$, and $i\in[k]$ because the $\delta$-path contains an alternating $a-c^{\prime}$ path and an alternating $a^{\prime}-c^{\prime}$ path. The edge $S_{i}$ is on both the path $P_{A}$ and the path of the $\delta$-path. We also know that $S_{i}$ is not in $P_{B}$, as the latter is disjoint from $P_{A}$. Since $B$ is on the cycle of the $\delta$-path, it is either in $Q_{j}$ for some $j\neq i$, or it is $S_{j}$ for some $j\neq i$. In both cases, either the $b-x_{B}$ or the $b^{\prime}-y_{B}$ subpath of $P_{B}$ has the property that it starts in $Q_{j}$ and does not contain $S_{j}$; we may assume by symmetry that it is the $b-x_{B}$ subpath. However, $x_{B}\in Q_{i}$ because of the edge $x_{A}x_{B}\in G^{\prime}$, so this subpath must contain $S_{i}$, a contradiction.
Next, consider the case $|V(M)\cap C|=2$. By Lemma 9, there is an $(A,B;C^{\prime})$-$\delta$-path such that $|V(M)\cap C^{\prime}|=1$. Using the first case for this $\delta$-path, we get that $A$ and $B$ are on the same alternating path in $M\Delta M_{0}$. ∎
@@
We define a subset $\mathcal{L}\subseteq\mathcal{K}$ the following way. A player $B$ is in $\mathcal{L}$, if for any $M_{0}$-alternating $A-B-C$ path for any $A\in\mathcal{K}$ and $C\in\mathcal{K}$, there exists an $(A,B;C)$-$\delta$-path or a $(C,B;A)$-$\delta$-path. Note that $\mathcal{K}^{*}\subseteq\mathcal{L}$ by Lemma 10.
@@
For $A,B\in\mathcal{L}$, we say that $(A,B)$ is a pair if there exists an $(A,B;C)$-$\delta$-path for some $C\in\mathcal{K}$. If $(A,B)$ is a pair and $M\in\mathcal{SC}$ is a matching such that $|V(M)\cap A|=1$, then by Lemma 9 there is an $(A,B;C^{\prime})$-$\delta$-path such that $|V(M)\cap B|=2$ and $|V(M)\cap C^{\prime}|=1$. By Proposition 11, $A$ and $B$ are on the same alternating path in $M\Delta M_{0}$.
Lemma 12.
Proof.
Suppose for contradiction that $(B,C)$ is not a pair. Let $M\in\mathcal{SC}$ be a matching such that $|V(M)\cap A|=1$, and let $P$ be the path in $M\Delta M_{0}$ that contains $A$. By Proposition 11, $B$ and $C$ are on $P$ and $|V(M)\cap B|=|V(M)\cap C|=2$. We may assume that $P$ is an $A-B-C-D$ path, where $D\in\mathcal{K}^{*}$ and $|V(M)\cap D|=1$. Let the order of the vertices of these players along the path be $a,a^{\prime},b,b^{\prime},c,c^{\prime},d,d^{\prime}$.
Since $(A,C)$ is a pair, there is an $(A,C;E)$-$\delta$-path $H$ for some $E\in\mathcal{K}^{*}$ by Lemma 9 such that $|V(M)\cap E|=1$. Furthermore, since $C\in\mathcal{L}$ and $(B,C)$ is not a pair, we obtain that there exists a $(D,C;B)$-$\delta$-path, by considering the definition of $\mathcal{L}$ for the $B-C-D$ subpath of $P$. Using Lemma 9, we can get a $(D,C;A)$-$\delta$-path $H^{\prime}$, since the two ends of the path of $M\Delta M_{0}$ containing $B$ are $A$ and $D$, so $A$ must be the tip of the $\delta$-path given by Lemma 9.
We now consider the graph $G^{\prime}$ obtained from $E\cup M_{0}$ by adding $d^{\prime}e^{\prime}$ and removing $A$, where $e^{\prime}$ is the tip of $H$. Since there is no $M_{0}$-alternating $a-a^{\prime}$ path in $G^{\prime}$, we can consider the GE-decomposition. The vertices $d^{\prime}$ and $e^{\prime}$ are in the same $Q_{i}$, since there are $M_{0}$-alternating paths to them from $a$ or $a^{\prime}$, and $d^{\prime}e^{\prime}$ is an edge in $G^{\prime}$. We also know that $i\in[k]$, because $H$ contains an alternating $a-e^{\prime}$ path and an alternating $a^{\prime}-e^{\prime}$ path.
In $H^{\prime}$, the $D-A$ path containing $C$ and the $D-A$ path disjoint from $C$ both contain $S_{i}$. This means that $S_{i}$ is on the path of $H^{\prime}$, and $C$ is in $Q_{i}$. However, in $H$, both the $A-E$ path containing $C$ and the $A-E$ path disjoint from $C$ contain $S_{i}$. Thus, $S_{i}$ is on the path of $H$, and therefore $C$ is outside of $Q_{i}$, a contradiction. ∎
@@
We are now ready to describe the structural properties that enable us to give a polynomial-time algorithm. Define $\mathcal{L}^{*}\subseteq\mathcal{L}$ the following way: a player $A\in\mathcal{L}$ is in $\mathcal{L}^{*}$ if for every $B,C\in\mathcal{L}$ such that $(A,B)$ and $(A,C)$ are pairs, $(B,C)$ is also a pair. It follows from Lemma 12 that $\mathcal{K}^{*}\subseteq\mathcal{L}^{*}$.
Since it can be decided in polynomial time whether an $(A,B;C)$-$\delta$-path exists for given $A,B,C\in\mathcal{K}$ (see Lemma 14), membership in $\mathcal{L}$ and $\mathcal{L}^{*}$ can be decided in polynomial time.
@@
We introduce an auxiliary graph $G^{*}$ whose vertex set is $\mathcal{L}^{*}$ and $(A,B)$ is an edge if and only if it is a pair. This graph can be constructed in polynomial time, and, by the definition of $\mathcal{L}^{*}$, every component of $G^{*}$ is a clique. Let $\mathcal{K}_{0}$ denote the players $A\in\mathcal{K}$ from which there is no $M_{0}$-alternating path to any other player in $\mathcal{K}$. Obviously, members of $\mathcal{K}_{0}$ are isolated vertices in $G^{*}$ (but there may be other isolated vertices).
Let $\mathcal{Q}$ denote the set of components of $G^{*}$. For a component $Q\in\mathcal{Q}$, let $V_{Q}\subseteq V$ denote the set of vertices of all players in $Q$. The existence of a polynomial-time algorithm for finding a member of $\mathcal{SC}$ or determining that $\mathcal{SC}=\emptyset$ follows from the following theorem.

### v2

@@ 2.2.1 Deciding the existence of a strong core matching
Our aim in this subsection is to prove that it can be decided in polynomial time if the strong core is nonempty, and if so, a matching in the strong core can be found in polynomial time. These results require a careful analysis of the structure of $M_{0}$-alternating paths, so the proof will feature several structural lemmas. We start by introducing some notation.
@@
Let $\mathcal{SC}$ denote the set of matchings in the strong core. We introduce some additional notation and definitions related to the structure of $M_{0}$-alternating paths.
$M_{0}$-alternating paths.
For three players $A$, $B$ and $C$, an $M_{0}$-alternating ABCA-B-C path is an $M_{0}$-alternating path that starts with the player edge $e_{A}$, ends with the player edge $e_{C}$, and also contains the player edge $e_{B}$.
$M_{0}$-alternating $\delta$-paths.
@@
An $M_{0}$-alternating $\delta$-path (or $\delta$-path for short) $H$ is an odd cycle and an $M_{0}$-alternating path with a common vertex $v$, where the $M_{0}$-edge adjacent to $v$ is on the path, and the cycle is $M_{0}$-alternating apart from $v$. The cycle of $H$ is denoted by $\mathrm{cyc}(H)$ and the path of $H$ is denoted by $\mathrm{pa}(H)$. An $(A,B;C)$-$\delta$-path for some players $A,B,C$ is an $M_{0}$-alternating $\delta$-path where $e_{A},e_{B}$ are in $\mathrm{cyc}(H)$ and $e_{C}$ is the end of the path $\mathrm{pa}(H)$. The vertex of $C$ at the end of the path is called the tip and the unique common vertex of $\mathrm{pa}(H)$ and $\mathrm{pa}(H)$ is called the connector. Note that the roles of $A$ and $B$ are exchangeable in the definition. See Figure 2 for an illustration.
@@
The sets of players $\mathcal{K},\mathcal{K}_{0},\mathcal{K}^{*},\mathcal{L},\mathcal{L}^{*}$.
We proceed to define some additional subsets and relations of players, that will be used extensively in the proofs.
  • Let $\mathcal{K}$ denote the set of players for whom there is no $M_{0}$-alternating cycle through the corresponding player edge.
  • Let $\mathcal{K}_{0}\subseteq\mathcal{K}$ denote the set of players $A\in\mathcal{K}$ from which there is no $M_{0}$-alternating path to any other player in $\mathcal{K}$.
  • We denote by $\mathcal{K}^{*}$ the set of players $A$ for whom $\exists M\in\mathcal{SC}$: $|V(M)\cap A|\leq 1$.
Of course, we cannot determine $\mathcal{K}^{*}$ without deciding whether the strong core is empty, so we will not use $\mathcal{K}^{*}$ in the algorithm, only in the proof of correctness. By Lemma 6 point a), we have $\mathcal{K}^{*}\subseteq\mathcal{K}$.
@@
  • We define a subset $\mathcal{L}\subseteq\mathcal{K}$ the following way. A player $B\in\mathcal{K}$ is in $\mathcal{L}$, if for any $M_{0}$-alternating ABCA-B-C path for any $A\in\mathcal{K}$ and $C\in\mathcal{K}$, there exists an $(A,B;C)$-$\delta$-path or a $(C,B;A)$-$\delta$-path.
@@
  • For $A,B\in\mathcal{L}$, we say that $(A,B)$ is a pair if there exists an $(A,B;C)$-$\delta$-path for some $C\in\mathcal{K}$.
@@
  • Define $\mathcal{L}^{*}\subseteq\mathcal{L}$ the following way: a player $A\in\mathcal{L}$ is in $\mathcal{L}^{*}$ if for every $B,C\in\mathcal{L}$ such that $(A,B)$ and $(A,C)$ are pairs, $(B,C)$ is also a pair.
The graph $G^{*}$.
@@
We introduce an auxiliary graph $G^{*}$ whose vertex set is $\mathcal{L}^{*}$ and $(A,B)$ is an edge if and only if it is a pair. We will show that this graph can be constructed in polynomial time. By the definition of $\mathcal{L}^{*}$, every connected component of $G^{*}$ is a clique. The members of $\mathcal{K}_{0}$ are isolated vertices in $G^{*}$ (but there may be other isolated vertices). Let $\mathcal{Q}$ denote the set of connected components of $G^{*}$. For a connected component $Q\in\mathcal{Q}$, let $V_{Q}\subseteq V$ denote the set of vertices of all players in $Q$.
We will illustrate these notions in an example below. But first, let us state our main theorem of the section about an alternative characterization of the strong core.
@@
An example.
We present two examples in Figure 3 in order to illustrate these new notions. The solid edges are in $M_{0}$, while the dashed edges are in $E$. In both examples, $\mathcal{K}=M_{0}$ because there is no $M_{0}$-alternating cycle, and $\mathcal{K}_{0}=\emptyset$. Also, $B\notin\mathcal{L}$ in both examples, since there is an ABXA-B-X alternating path, but there is neither an $(A,B;X)$-$\delta$-path nor an $(X,B;A)$-$\delta$-path. It can be checked that the remaining players are in $\mathcal{L}$ in both examples.
Figure 3: Solid edges are in $M_{0}$, dashed edges are in $E$. The example on the right is obtained from the example on the left by adding $ab^{\prime}$. The strong core is empty on the left, while bold edges on the right form a matching in the strong core.
In the left example, the pairs are $(A,D),(A,X),(C,D),(C,X),(D,X)$, which means that $\mathcal{L}^{*}=\{A,C,Y\}$, and each are singletons in $G^{*}$. It can be checked that there is no matching that covers each of $A,C,X$ at least once and the remaining players twice, so $\mathcal{SC}=\emptyset$.
In the right example, $(A,C)$ is also a pair, so $\mathcal{L}^{*}=\{A,C,D,X,Y\}$, where $Y$ is a singleton in $G^{*}$, while $A,C,D,X$ form a clique. By Theorem 8, $M\in\mathcal{SC}$ if and only if it covers both vertices of $B$ and it leaves at most one vertex of $Y$ and at most one vertex of $A\cup C\cup D\cup X$ exposed. One such matching $M$ is indicated in bold; the dotted path is the single alternating path in $M\Delta M_{0}$.
We will see at the end of the section that Theorem 8 not only allows us to decide in polynomial time whether the strong core is empty, it also enables us to minimize a linear cost function on the strong core in polynomial time. Next, we proceed to prove the theorem.
Proving Theorem 8
We start with the key lemma, which states that in many cases, we can either “unwind” alternating paths or create a $\delta$-path. In order to be able to refer to the objects in the lemma in various contexts later in the proof, we use subscript or superscript $L$ to denote the objects in the conditions of the lemma. See Figure 4 for an illustration.
Figure 4: Illustration of the Unwinding Lemma. Top: the $M_{0}$-alternating $A_{L}-B_{L}-C_{L}$ path $P^{L}$, and its subpath $P^{L}_{bc^{\prime}}$. Middle: The $M_{0}$-alternating path $P^{L}_{A_{L}}$ through the player edge $e_{A_{L}}=aa^{\prime}$, and its subpaths $P^{L}_{a}$ and $P^{L}_{a^{\prime}}$. Bottom: in this example, the lemma gives an $a^{\prime}-b$ path $P_{a^{\prime}b}$ such that $P=P_{a^{\prime}b}\cup P^{L}_{bc^{\prime}}$ is an $M_{0}$-alternating path, and $P^{L}_{a^{\prime}}\cup P$ contains an $(A_{L},B_{L};C_{L})$-$\delta$-path $H$ with $P\subseteq H$.
Lemma 9 (Unwinding Lemma).
Let $A_{L}\in\mathcal{K}$, and let $P^{L}_{A_{L}}$ be an $M_{0}$-alternating path through $e_{A_{L}}$. Furthermore, let $P^{L}$ be an $A_{L}-C_{L}$ $M_{0}$-alternating path that ends in $c^{\prime}\in C_{L}$ for some player $C_{L}$, and let $A_{L}\neq B_{L}=\{b,b^{\prime}\}$ be a player whose edge $bb^{\prime}$ is on this path, where $b^{\prime}$ is the vertex closer to $c^{\prime}$. We denote by $P^{L}_{bc^{\prime}}$ the part of $P^{L}$ from $b$ to $c^{\prime}$.
Under these conditions, there exist $u\in A_{L}$ and an $M_{0}$-alternating ubu-b path $P_{ub}\subseteq P^{L}_{A_{L}}\cup P^{L}$ such that $P=P_{ub}\cup P^{L}_{bc^{\prime}}$ is an $M_{0}$-alternating path, and either
  • the subpath $P^{L}_{u}$ of $P^{L}_{A_{L}}$ starting in $u$ is disjoint from $P$ (apart from $u$), i.e., $P^{L}_{u}\cup P$ is an $M_{0}$-alternating path, or
  • $P^{L}_{u}\cup P$ contains an $(A_{L},B_{L};C_{L})$-$\delta$-path $H$ with $P\subseteq H$.